OpenStudy (anonymous):

Please help!!! What is the least common multiple for: 12x^2y^3 and 14x^3y^2 Thanks!

OpenStudy (anonymous):

Well, let us first factorize both as much as we can: 12x^2y^3 = 2*2*3*x^2*y^3 14x^3y^2 = 2*7*x^3*y^2 Common factors are 2*x^2*y^2, so we will only use these once. Therefore lcm = 2*3*y*7*x*2*x^2*y^2 = 84x^3y^3